# Ego-VPA: Egocentric Video Understanding with Parameter-efficient Adaptation

### v1

@@ 1 Introduction
Figure 1: Ego-VPA leverages context-aware prompts to achieve parameter-efficient adaptation for egocentric videos. (Left) Performance vs tunable parameters; (Right) Cross modality prompt-tuning in Ego-VPA.
@@
Table 1: The zero-shot - fine-tuned performance gap exists in both CLIP-based VFMs [25, 42] and Ego-VFMs [49] (mAP on Charades-Ego). denotes that only the prompts/adapters are fine-tuned on Charades-Ego.
We consider instead the lightweight adaptation of VFMs pretrained on egocentric videos [20, 29, 49] (Ego-VFMs), for which we propose efficient prompt-tuning techniques. Since no prior work investigated this problem, we start by proposing several baselines, including adding learnable prompts only to the text or the video encoder. Since single-modality prompts fail to capture connections between the text and spatial-temporal context, they unsurprisingly lead to marginal improvements over zero-shot VFM performance. We then show that better results can be obtained by prompt-tuning both the video and text encoders, using the recent VoP approach [12], which uses an additional module, a bi-directional LSTM, to connect visual prompts across frames, according to context. While achieving good performance, this introduces a large parameter overhead, compromising the lightweight nature of the adaptation (See trainable parameters in Table 2).
@@
Overall, we make three contributions to the efficient adaptation to egocentric video understanding. First, we show that prompt-tuning video encoders based on IFMs are suboptimal for egocentric video tasks due to the inherent domain gap. Second, we propose several baselines that prompt-tune the existing Ego-VFMs and show that these baselines are not effective and efficient. Finally, we propose a novel and efficient prompt-tuning approach, Ego-VPA, that utilizes a local subspace approximation with shared basis prompts for cross-modal prompt synthesis, enabling context reasoning across frames and modalities.
@@ 3.1 VFM Preliminaries
Video encoder $\phi_{vid}$. Given a video $V$ with $T$ frames, each frame ${\bf V}^{f}$ is mapped into patch embeddings $\{{\bf z}_{(f,p)}^{(0)}\}_{p=1}^{N_{p}}$, where $N_{p}$ is the number of patches. A [CLS] token ${\bf z}_{[CLS]}^{(0)}\in\mathbb{R}^{d_{vid}}$ is prepended to the input sequence, i.e. ${\bf Z}^{(0)}_{vid}=({\bf z}_{[CLS]}^{(0)},{\bf z}_{(1,1)}^{(0)},...,{\bf z}_{(T,N_{p})}^{(0)})^{T}\in\mathbb{R}^{(N_{p}T+1)\times d_{vid}}$. The video feature $\phi_{vid}(V)={\bf z}^{(L)}_{[CLS]}$ is then extracted with the video encoder $\phi_{vid}$ with transformer mapping of Eq.(1). Since the $(N_{p}T+1)$ attention computations of each patch induce large memory overhead, there is a need to trade-off between space and time. VFMs typically use the TimeSformer [3] architecture, which utilizes two types of attention: a spatial attention block that only attends to features from the same time/frame (i.e. $({\bf z}_{[CLS]},{\bf z}_{(f,1)},...,{\bf z}_{(f,N_{p})})$); and a time attention block that only attends to features from the same location (i.e., $({\bf z}_{[CLS]},{\bf z}_{(1,p)},...,{\bf z}_{(T,p)})$). The two attention blocks are interleaved to construct each transformer block of Eq. (1), reducing the attention computations per patch to $(N_{p}+T+2)$. This enables increasing the sampling rate of each video (e.g., from 4 frames in [2] to 16 frames).
@@ 4 Ego-VFM Prompt-tuning Baselines
Video prompt-tuning (VPT). Given a video $V$, the video embedding of ${\bf Z}_{vid}^{(0)}\in\mathbb{R}^{(N_{p}T+1)\times d_{vid}}$ is extracted as described in section 3.1. To prompt-tune the video encoder, the visual prompts ${\bf P}_{v}=({\bf p}_{v,1},...,{\bf p}_{v,M_{v}})\in\mathbb{R}^{d_{vid}\times M_{v}}$ are prepended to the input sequence, i.e. $\tilde{\bf Z}_{vid}^{(0)}=({\bf z}^{(0)}_{[CLS]},{\bf P}_{v},{\bf z}_{(1,1)}^{(0)},...,{\bf z}_{(T,N_{p})}^{(0)})^{T}$. Note that we apply prompt-tuning only to the spatial attention blocks of the TimeSformer [3] as we found that prompt-tuning both blocks has no additional gain (See Appendix). Inspired by [13], beyond introducing visual prompts at the input layer, we further prepend layer-specific prompts ${\bf P}_{v}^{(l)}$ to every transformer layer. In the following, we omit the layer $l$ superscript of these deep prompts, for brevity. The output of VPT is then written as $\phi_{vid}^{\prime}(V)=\tilde{\bf z}^{(L)}_{[CLS]}$.
@@ 5 Ego-VPA
Our experiments (see Section 6.2) show that $VoP^{F+C}$ prompting significantly improves the performance of the Ego-VFM on new egocentric datasets. However, this method still falls short in two aspects. First, the CMM module, which is itself a bi-directional LSTM network, still requires a substantial number of parameters. Second, there is no connection between text and visual prompts, limiting knowledge transfer across modalities. We next propose a novel prompt-tuning technique for Ego-VFMs such as LaViLa, denoted as Ego-VPA, which achieves context fusion across frames (like CMM) and modalities in an extremely lightweight fashion. This is implemented by the proposed prompt synthesis scheme, using a shared prompt basis. Note that we keep the inter-frame attention layers (last $L-K$ layers) identical to $VoP^{F+C}$ for fair comparison.
@@ 5.1 Video Prompt Synthesis
where $\mathbf{F}\in\mathbb{R}^{d_{f}\times B}$ contains the $B$ basis prompts as columns and $||\alpha||_{0}$ is the 0-norm (number of non-zero elements) of $\alpha$. We refer to this as the local reconstruction subspace for $h_{vid}({\bf z}_{f})$ and denote the set of vectors
@@
The vector $\alpha_{f}^{*}$ has a closed form solution due to the orthogonality of the basis $\cal F$. For any matrix $\mathbf{G}$ containing $k$ columns of $\bf F$, the minimizer of $\|h-\mathbf{G}\alpha\|_{2}$ is
$\alpha=({\bf G}^{T}{\bf G})^{-1}{\bf G}^{T}h={\bf G}^{T}h_{vid}({\bf z}_{f}),$ (5)
@@
$\|h-\mathbf{G}\alpha\|_{2}=(h-\mathbf{G}\alpha)^{T}(h-\mathbf{G}\alpha)~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}=\|h\|_{2}-2h^{T}{\bf G}\alpha+\|\alpha\|_{2}=\|h\|_{2}-\|\alpha\|_{2}.$ (6)
@@ 6 Experiments
In this section, we validate the efficiency and effectiveness of Ego-VPA by comparing it with SOTA methods and ablating on different components.
6.1 Experimental Setup
Datasets. We adopt LaViLa [49] as the Ego-VFM, which is pretrained on Ego4D [8], containing 4M video-text pairs for egocentric videos. Both the proposed method and baselines are evaluated on Charades-Ego [34], EGTEA [19], and EPIC-Kitchens-100 [6], which are widely used in egocentric video research. Charades-Ego [34] is a fine-grained action classification dataset containing $33,114$ trimmed action segments for training, spanning across 157 classes. While both egocentric and exocentric views are provided, we only train and evaluate egocentric videos using the official splits, as in  [20, 49]. EGTEA [19] is an egocentric cooking video dataset, containing $10,321$ action instances from 106 fine-grained classes. We only use the video data and follow the same protocol for training and evaluation as in [49]. EPIC-Kitchens-100 [6] is an egocentric cooking video dataset of $100$ hours, containing $67,217$/$9,668$ clips for training/validation. We evaluate the multi-instance retrieval (MIR) task to test the generalization of Ego-VPA, which is a text-to-video (T->V) and video-to-text (V->T) retrieval task.
@@ 6.2 Comparisons to SOTA Prompt-tuning Methods
Table 2 presents the result on Charades-Ego and EGTEA, as a function of the tunable model parameters. The poor zero-shot performance shows that current Ego-VFMs, like LaViLa, are somewhat overfitted to the Ego4D dataset. Full fine-tuning significantly improves performance, but requires optimization of the entire VFM, which is inefficient. Prompt-tuning methods require less parameter optimization. TPT [51] and VPT [13] extend the input sequence with learnable prompts for the text and video encoders, respectively. These methods are weaker than VoP [12], where prompt-tuning is performed for both encoders. VoPF+C, a variant of VoP using a CMM module and frame-aware attention layers, is the best-performing baseline. However, the introduction of the CMM module significantly increases the parameter counts, requiring around $10\%$ of the model size for the adaptation. The proposed Ego-VPA consistently outperforms all other methods on both datasets, even beating full fine-tuning, with only $0.84\%$ trainable parameters. While the adaptation to EGTEA produces much stronger results, indicating that the domain gap to Ego4D is smaller, the zero-shot performance of LaViLa is still quite weak, reinforcing the importance of efficient adaptation methods.
@@ 6.3 Ablation Studies
Context modeling module. We first compare the proposed prompt synthesis (PS) with the CMM module of $VoP^{F+C}$ [12]. As shown in Table 3, despite using much fewer trainable parameters, video-only prompt synthesis (m3) outperforms $VoP^{F+C}$ (m1). When cross-modal knowledge transfer is enabled with cross-modal prompt synthesis (m5), the performance further improves by a significant margin. This validates the effectiveness of context modeling across frames and modalities.
@@
Number of frames. We validate the robustness of Ego-VPA by training with different numbers of frames on both Charades-Ego and EGTEA. As shown in Table 4, the performance of all methods increases with the number of frames, indicating that temporal resolution is an important factor for video understanding. Ego-VPA consistently improves over the VoPF+C and the full fine-tuning performances, showing that it is robust and effective across time resolutions.
@@ 6.4 Generalization to Retrieval Tasks
We further evaluate multi-instance retrieval tasks [6] on Epic-Kitchens-100. Figure 6 shows that Ego-VPA performs on par with VoPF+C, which has $10\%$ more trainable parameters. Compared to vanilla VoP, which has a similar number of parameters, Ego-VPA has clearly better performance. This shows that Ego-VPA is more parameter-efficient and effective across different egocentric video tasks.
@@ 7 Conclusions
We propose Ego-VPA, a novel parameter-efficient adaptation method for Ego-VFMs. Atop a frozen Ego-VFM, we sparsely approximate projected video frame/text features with a shared set of basis prompts and synthesize video/text prompts accordingly. This is shown to enhance context fusion across frames and cross-modal transfer, reaching better visual-language alignments. Through extensive experiments, we demonstrate that Ego-VPA is both efficient and effective, outperforming state-of-the-art methods with much fewer learnable parameters.

### v2

@@ 1 Introduction
Figure 1: Ego-VPA leverages context-aware prompts to achieve parameter-efficient adaptation for egocentric videos. (Left) Performance vs tunable parameters; (Right) Cross modality prompt-tuning in Ego-VPA where the VFM is frozen.
@@
Table 1: The zero-shot - fine-tuned performance gap (mAP on Charades-Ego) exists in both CLIP-based VFMs [26, 43] and Ego-VFMs [51] . denotes that only the prompts/adapters are fine-tuned on Charades-Ego.
In this work, we investigate efficient prompt tuning based frameworks to address light-weight adaptation of VFMs, in particular Ego-VFMs [21, 30, 51] which are pretrained on egocentric videos. Since no prior work investigated this problem, we begin with introducing several baselines, including adding learnable prompts only to the text or the video encoder. Single-modality prompts unsurprisingly lead to marginal improvements over zero-shot VFM performance as they fail to capture connections between the text and spatial-temporal context. We then show that improved results can be obtained by prompt-tuning both the video and text encoders, using the recent VoP approach [13], which uses an additional module, a bi-directional LSTM, to connect visual prompts across frames, according to context. While achieving good performance, this introduces a large parameter overhead, compromising the lightweight nature of the adaptation (See trainable parameters in Table 2).
@@
Overall, we make three contributions to the efficient adaptation for egocentric video understanding. First, we show that prompt-tuning video encoders based on IFMs are suboptimal for egocentric video tasks due to the inherent domain gap. Second, we propose several baselines that prompt-tune the existing Ego-VFMs and show that these baselines are not effective and efficient. Finally, we propose a novel and efficient prompt-tuning approach, Ego-VPA, that utilizes a local subspace approximation with shared basis prompts for cross-modal prompt synthesis, enabling context reasoning across frames and modalities.
@@ 3.1 VFM Preliminaries
Video encoder $\phi_{vid}$. Given a video $V$ with $T$ frames, each frame ${\bf V}^{f}$ is mapped into patch embeddings $\{{\bf z}_{(f,p)}^{(0)}\}_{p=1}^{N_{p}}$, where $N_{p}$ is the number of patches. A [CLS] token ${\bf z}_{[CLS]}^{(0)}\in\mathbb{R}^{d_{vid}}$ is prepended to the input sequence, i.e. ${\bf Z}^{(0)}_{vid}=({\bf z}_{[CLS]}^{(0)},{\bf z}_{(1,1)}^{(0)},...,{\bf z}_{(T,N_{p})}^{(0)})^{T}\in\mathbb{R}^{(N_{p}T+1)\times d_{vid}}$. The video feature $\phi_{vid}(V)={\bf z}^{(L)}_{[CLS]}$ is then extracted with the video encoder $\phi_{vid}$ with transformer mapping of Eq. (1). Since the $(N_{p}T+1)$ attention computations of each patch induce large memory overhead, there is a need to trade-off between space and time. VFMs typically use the TimeSformer [3] architecture, which utilizes two types of attention: a spatial attention block that only attends to features from the same time/frame (i.e. $({\bf z}_{[CLS]},{\bf z}_{(f,1)},...,{\bf z}_{(f,N_{p})})$); and a time attention block that only attends to features from the same location (i.e., $({\bf z}_{[CLS]},{\bf z}_{(1,p)},...,{\bf z}_{(T,p)})$). The two attention blocks are interleaved to construct each transformer block of Eq. (1), reducing the attention computations per patch to $(N_{p}+T+2)$. This enables increasing the sampling rate of each video (e.g., from 4 frames in [2] to 16 frames). Note that this is especially important for egocentric videos as camera tends to move faster.
@@ 4 Ego-VFM Prompt-tuning Baselines
Video prompt-tuning (VPT). Given a video $V$, the video embedding of ${\bf Z}_{vid}^{(0)}\in\mathbb{R}^{(N_{p}T+1)\times d_{vid}}$ is extracted as described in section 3.1. To prompt-tune the video encoder, the visual prompts ${\bf P}_{v}=({\bf p}_{v,1},...,{\bf p}_{v,M_{v}})\in\mathbb{R}^{d_{vid}\times M_{v}}$ are prepended to the input sequence, i.e. $\tilde{\bf Z}_{vid}^{(0)}=({\bf z}^{(0)}_{[CLS]},{\bf P}_{v},{\bf z}_{(1,1)}^{(0)},...,{\bf z}_{(T,N_{p})}^{(0)})^{T}$. Note that we apply prompt tuning only to the spatial attention blocks of the TimeSformer [3] as we found that prompt-tuning both blocks has no additional gain (See Appendix). Inspired by [14], beyond introducing visual prompts at the input layer, we further prepend layer-specific prompts ${\bf P}_{v}^{(l)}$ to every transformer layer. In the following, we omit the layer $l$ superscript of these deep prompts, for brevity. The output of VPT is then written as $\phi_{vid}^{\prime}(V)=\tilde{\bf z}^{(L)}_{[CLS]}$.
@@ 5 Ego-VPA
Our experiments (see Section 6.2) show that $VoP^{F+C}$ prompting significantly improves the performance of the Ego-VFM on new egocentric datasets. However, this method still falls short in two aspects. First, the CMM module, which is itself a bi-directional LSTM network, still requires a substantial number of model parameters. Second, there is no connection between text and visual prompts, limiting knowledge transfer across modalities. We next propose a novel prompt-tuning technique for Ego-VFMs such as LaViLa, denoted as Ego-VPA, which achieves context fusion across frames (like CMM) and modalities in an extremely lightweight fashion. This is implemented by the proposed prompt synthesis scheme, using a shared prompt basis. Note that we keep the inter-frame attention layers (last $L-K$ layers) identical to $VoP^{F+C}$ for fair comparison.
@@ 5.1 Video Prompt Synthesis
where $\mathbf{F}\in\mathbb{R}^{d_{f}\times B}$ contains the $B$ basis prompts in $\cal F$ as columns and $||\alpha||_{0}$ is the 0-norm (number of non-zero elements) of $\alpha$. We refer to this as the local reconstruction subspace for $h_{vid}({\bf z}_{f})$ and denote the set of vectors
@@
The vector $\alpha_{f}^{*}$ has a closed form solution due to the orthogonality of the basis $\cal F$. For any matrix $\mathbf{G}$ containing $k$ columns of $\bf F$, the minimizer of $\|{\bf h}-\mathbf{G}\alpha\|_{2}$ is
$\alpha=({\bf G}^{T}{\bf G})^{-1}{\bf G}^{T}{\bf h}={\bf G}^{T}h_{vid}({\bf z}_{f}),$ (5)
@@
$\|{\bf h}-\mathbf{G}\alpha\|_{2}=({\bf h}-\mathbf{G}\alpha)^{T}({\bf h}-\mathbf{G}\alpha)~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}=\|{\bf h}\|_{2}-2{\bf h}^{T}{\bf G}\alpha+\|\alpha\|_{2}=\|{\bf h}\|_{2}-\|\alpha\|_{2}.$ (6)
@@ 6 Experiments
In this section, we validate the efficiency and effectiveness of Ego-VPA by comparing it with SOTA methods and ablating on different components. More results in the appendix.
6.1 Experimental Setup
Datasets. We adopt LaViLa [51] as the Ego-VFM, which is pretrained on Ego4D [9], containing 4M video-text pairs for egocentric videos. Both the proposed method and baselines are evaluated on Charades-Ego [35], EGTEA [20], and EPIC-Kitchens-100 [7], which are widely used in egocentric video research. Charades-Ego [35] is a fine-grained action classification dataset containing $33,114$ trimmed action segments for training, spanning across 157 classes. While both egocentric and exocentric views are provided, we only train and evaluate egocentric videos using the official splits, as in  [21, 51]. EGTEA [20] is an egocentric cooking video dataset, containing $10,321$ action instances from 106 fine-grained classes. We only use the video data and follow the same protocol for training and evaluation as in [51]. EPIC-Kitchens-100 [7] is an egocentric cooking video dataset of $100$ hours, containing $67,217$/$9,668$ clips for training/validation. We evaluate the multi-instance retrieval (MIR) task to test the generalization of Ego-VPA, which contains a text-to-video (T->V) and video-to-text (V->T) retrieval task.
@@ 6.2 Comparisons to SOTA Prompt-tuning Methods
Table 2 presents the result on Charades-Ego and EGTEA, as a function of the tunable model parameters. The poor zero-shot performance shows that current Ego-VFMs, like LaViLa, are somewhat overfitted to the Ego4D dataset. Full fine-tuning significantly improves performance, but requires optimization of the entire VFM, which is inefficient. Prompt-tuning methods require less parameter optimization. Bias [5, 48] only fine-tunes the bias terms in the model but not the weights, and thus has limited adaptation capacity. TPT [53] and VPT [14] extend the input sequence with learnable prompts for the text and video encoders, respectively. These methods are weaker than VoP [13], where prompt-tuning is performed for both encoders. VoPF+C, a variant of VoP using a CMM module and frame-aware attention layers, is the best-performing baseline. However, the introduction of the CMM module significantly increases the parameter counts, requiring around $10\%$ of the model size for the adaptation. The proposed Ego-VPA consistently outperforms all other methods on both datasets, even beating full fine-tuning, with only $0.84\%$ trainable parameters. While the adaptation to EGTEA produces much stronger results, indicating that the domain gap to Ego4D is smaller, the zero-shot performance of LaViLa is still quite weak, reinforcing the importance of efficient adaptation methods.
@@ 6.3 Ablation Studies
Context modeling module. We first compare the proposed prompt synthesis (PS) with the CMM module of $VoP^{F+C}$ [13]. As shown in Table 3, despite using much fewer trainable parameters, video-only prompt synthesis (m3) outperforms the CMM in $VoP^{F+C}$ (m1). When cross-modal knowledge transfer is enabled with cross-modal prompt synthesis (m5), the performance further improves largely. This validates the effectiveness of context modeling across frames and modalities.
@@
Number of frames. We validate the robustness of Ego-VPA by training with different numbers of frames on both Charades-Ego and EGTEA. As shown in Table 4, the performance of all methods increases with the number of frames, indicating that temporal resolution is an important factor for video understanding. Ego-VPA consistently improves over $VoP^{F+C}$ and the full fine-tuning performances, showing that it is robust and effective across time resolutions.
@@ 6.4 Generalization to Retrieval Tasks
We further evaluate multi-instance retrieval tasks [7] on Epic-Kitchens-100. Figure 6 shows that Ego-VPA performs on par with $VoP^{F+C}$, which has $10\%$ more trainable parameters. Compared to vanilla VoP, which has a similar number of parameters, Ego-VPA has clearly better performance. This shows that Ego-VPA is more parameter-efficient and effective across different egocentric video tasks.
@@ 7 Conclusions
We propose Ego-VPA, a novel parameter-efficient adaptation method for Ego-VFMs. Atop a frozen Ego-VFM, we sparsely approximate projected video frame/text features with a shared prompt basis and synthesize video/text prompts accordingly. This is shown to enhance context fusion across frames and cross-modal transfer, achieving improved visual-language alignments. Through extensive experiments, we show that Ego-VPA is both efficient and effective, outperforming SOTA methods with much fewer learnable parameters.
Acknowledgment This work was partially funded by NSF awards IIS-2303153, and a gift from Qualcomm.